In order to conduct a hypothesis test for the population proportion, you sample 290 observations that result in 87 successes. (You may find it useful to reference the appropriate table: z table or t table)
H0: p = 0.36; HA: p ≠ 0.36.
1. Calculate the value of the test statistic. (Negative value should be indicated by a minus sign. Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)

2. Find the p-value.
A. p-value > or equal to 0.10
B. p-value < 0.01
C. 0.01 < or equal to p-value < 0.025
D. 0.025 < or equal to p-value < 0.05
E. 0.05 < or equal to p-value < 0.10
3. At the 0.01 significance level, What is the conclusion?
A. Reject H0 since the p-value is greater than significance level.
B. Reject H0 since the p-value is smaller than significance level.
C. Do not reject H0 since the p-value is greater than significance level.
D. Do not reject H0 since the p-value is smaller than significance level.
4. Interpret the results at α = 0.01.
A. We conclude that the population mean differs from 0.36.
B. We cannot conclude that the population mean differs from 0.36.
C. We conclude the population proportion differs from 0.36.
D. We cannot conclude that the population proportion differs from 0.36.

Answer by Boreal(15235)

test stat is a z=(phat-p)/sqrt(p*(1-p)/n); phat=87/290=0.3; sqrt term is sqrt(0.36*0.64/290)=0.0282
z=-0.06/0.0282
=-2.13
p-value is curve on outside of +2.1286 and -2.1286, since 2 way test
that is 0.03228
Fail to reject at the 0.01 level since p-value is greater than the signficance level. We can conclude that the population proportion differs from 0.36.